OpenStudy (anonymous):

Determine the seating capacity of an auditorium with 18 rows of seats if there are 20 seats in the first row, 24 seats in the second row, and 28 seats in the third row and so on.

OpenStudy (anonymous):

lol that's a weird auditorium. quarter circle.

OpenStudy (anonymous):

it has 20*18+4(1+2+3+...+17) seats

OpenStudy (anonymous):

= 20*18 +4(17*16)/2

OpenStudy (anonymous):

i'm sorry, do you think you can explain it? i just don't understand what you have to do.

OpenStudy (anonymous):

the capacity of the auditorium is 1102

OpenStudy (anonymous):

it's 972. sorry first answer is incorect

OpenStudy (anonymous):

the first row has 20 seats. the second row has 24 seats and the third row has 28 seats and so on. so it is 20+24+28+.... 20+20+20+...+4+8+12... 20*18+4(1+2+3+...+17) = 20*18 +4(17*18/2)

OpenStudy (anonymous):

= 972

OpenStudy (anonymous):

since 1+2+3+4+....+ n-1 +n = n(n+1)/2